OpenStudy (anonymous):

how would I find the nth term in the sequence 1 0 25 0 625. I think the sin function has to be involved.

OpenStudy (anonymous):

Do you need to write a function to solve this? I would just notice what is given and you can infer the nth term. Your sequence is based on f(n)=5^n, where n is an integer greater than or equal to 0. n=0 --> 5^n = 5^0 = 1 n=1 --> 0 n=2 --> 5^n = 5^2 = 25 n=3 --> 0 n=4 --> 5^n = 5^4 = 625 When n is even, your sequence returns the value of the function. When n is odd, your function returns 0. So if you wanted to find the nth term, whatever that term is, if you define n as an odd number, the value of the nth term is 0. If you define n as an even number, the value of the nth term is 5^n

OpenStudy (anonymous):

why do odd numbers return to 0? Isn't 5^1 =5 and 5^3=125?

OpenStudy (anonymous):

Yes, that's what I'm unsure of. I'm not sure if sitting on it longer would help me figure out a function for it specifically, but you are perfectly correct. For whatever reason, your function here returns 0 for all odd-numbered powers.

OpenStudy (anonymous):

Maybe piece-wise actually.... Hang on that might work here! Give me a sec.

OpenStudy (anonymous):

I mean, I don't think this is needed, but I guess it technically would work here.